Could I get some help with graphing this equation: 4x - 3y = 12. Thank You
Answer by AnlytcPhil(1810)

Could I get some help with graphing this equation: 4x - 3y = 12? 
Thank You

1.  Find the x-coordinate of the x-intercept by substituting 0 for y and
    solving for x.  The y-coordinate of the x-intercept is 0.

      4x - 3y = 12
    4x + 3(0) = 12
           4x = 12
            x = 12/4
            x = 3

    So the x-intercept is (3,0)

2.  Find the y-coordinate of the y-intercept by substituting 0 for x and
    solving for y.  The x-coordinate of the y-intercept is 0.

     4x - 3y = 12
   4(0) - 3y = 12
         -3y = 12
           y = -4
 
    So the x-intercept is (0,-4)
   
3. Plot these two points: (3,0) is on the x-axis 3 units right of the
   origin and (0,-4) is on the y-axis 4 units below the origin. 

           
    
Then draw a line through those two points:

          

Edwin